Would you please help me with this problem
The questions is to simplify and combine like radicals. All variables represent positive numbers.
sqrt2 - sqrt8
My results
sqrt2 = 1*2
sqrt8 = 2*2*2
answer -sqrt4 ( I was told that this is the wrong answer)
Please help!!!

Answer by Fombitz(13823)

Let's start from your results
sqrt2 = 1*2
sqrt8 = 2*2*2
Your right hand side should still include the square root.
sqrt2 = sqrt(1*2)
sqrt8 = sqrt(2*2*2)
Then
sqrt2 = sqrt(2)
sqrt8 = sqrt(2*2)*sqrt(2)
and you can substitute since
sqrt(2*2) = 2
Your equations then become
sqrt2 = sqrt(2)
sqrt8 = 2*sqrt(2)
Now your subtraction can take place.
sqrt2 - sqrt8 = sqrt2 - 2*sqrt(2)
sqrt2 - sqrt8 = -sqrt(2)